OpenStudy (anonymous):

Express the complex number in trigonometric form. -2i

ganeshie8 (ganeshie8):

\(a + bi = r(\cos \theta + i \sin \theta)\)

ganeshie8 (ganeshie8):

maybe plot the complex number in plane, u wud see what its angle and magnitude are

ganeshie8 (ganeshie8):